OpenStudy (arbyj):

How do you turn 4x-y=2 into slope-intercept form

OpenStudy (anonymous):

y = mx + b is slope-intercept form. so you have to get 'y' by itself on one side of the = sign

OpenStudy (anonymous):

4x-y=2 -y=-4x+2 y=4x-2

OpenStudy (arbyj):

thank you :)